I'm really lost on how to do these.
Depreciation,solve the literal equation for the indicated variable.
D = c- S
______ (for s)
N


Answer by stanbon(75887)

Solve D = (c- S)/N for S.
Multiply both sides by N to get:
DN=c-S
Add S to both sides and subtract DN from both sides to get:
S=c-DN
Cheers,
Stan H.